please help me find these coordinates.
the coordinate of point T are given. the midpoint of ST has coordinate (5,-8). find the coordinates of points S.
1. T(0,4)
2. T(10,18)

Answer by stanbon(75887)

the coordinate of point T are given. the midpoint of ST has coordinate (5,-8). find the coordinates of points S.
You have to know that the coordinates of the midpoint are the average of
the x-values of the end points and the y-values of the end points.
1. T(0,4)
Let S have coordinates (x,y)
Then (x+0)/2 = 5; x=10
And (y+4)/2 = -8; y=-20
-----------------

2. T(10,18)
Let S have coordinates (x,y)
Then (x+10)/2 = 5; x=0
And (y+18)/2 =-8; y=-34
Cheers,
Stan H.